OpenStudy (anonymous):

7. Which is the direct linear variation equation for the relationship? y varies directly with x and y = 12 when x = 4. (Points : 4) y = 2x + 4 y = 3x y = x + 8 y = x – 8

OpenStudy (anonymous):

y varies directly with x is written as: \[y \alpha x\] To make it into an equation we introduce the proportionality constant, k, y=kx So at x=4 and y=12, we substitute into the equation to get 12=k*4 ie k=12/4=3 So the answer is: y=3x.